OpenStudy (anonymous):

Solve the following system of linear equations: 16x-16y+24z=120 -13x +13y -20z =-99 51x -51y +77z =384 I know there are infinitely many solutions with 1 arbitrary parameter(which is y=y) Im having trouble with x and z. Help?

OpenStudy (anonymous):

treat y like any constant and find x and z like functions of y

OpenStudy (anonymous):

still not getting the answer. if i put a 1 in for y, i solve to x=13/2-3/2z and z=13/3-2/3x which is wrong, where am i going wrong?

OpenStudy (anonymous):

y = x-3 z = 3

OpenStudy (anonymous):

put y = p for example then rearange equations like this: 16x+24z=120+16p -13x-20z=-99-13p 51x+77z=384+51p so now you have 3 equations qith 2 uknown. Choose any 2 of them and solve them to find x and z

OpenStudy (anonymous):

ok, I will try that. Thank you for your help!

OpenStudy (anonymous):

yw